# SEW: A full-spectrum linear fitting with stellar population synthesis method Based on “Equivalent Widths spectrum”

### v1

@@ 1 Introduction
In Galaxy spectra analysis, two primary methods are commonly employed: the full-spectrum fitting with stellar population synthesis (SPS) method and the absorption line characterisation method.
In SPS, the observed spectrum is modelled as a linear combination of SSPs, each with specific age and metallicity. These SSPs are integrated over an initial mass function (IMF) and stellar evolution model. The fitting process adjusts the weights of SSPs, along with other parameters, to minimise the difference between the model and data:
$F(\lambda)$ $=F_{\text{init}}(\lambda)*P(\lambda),$ (1a)
$F_{\text{init}}(\lambda)$ $=\sum_{j}x_{j}f_{\text{temp}}(j)(\lambda),$ (1b)
@@
Here, $f_{\text{temp}}(j)(\lambda)$ represents SSP and emission line templates, $x_{j}$ is their weights, and $P(\lambda)$ includes the global dust attenuation curve $A(\lambda)$ and a correction function $q(\lambda)$ for systematic bias. This approach allows for detailed decomposition of stellar populations and estimation of various galaxy properties. A number of spectral synthesis fitting codes are now publicly available and are commonly adopted, such as pPXF developed by Cappellari & Emsellem (2004); Cappellari (2017), STARLIGHT by Cid Fernandes et al. (2005), STECKMAP by Ocvirk et al. (2006a, b), and VESPA by Tojeiro et al. (2007). However, the accuracy of the fitting results is significantly influenced by several interrelated factors, including the choice of the theoretical dust attenuation curve, the precision of spectral calibrations, and the inherent degeneracy among age, metallicity, and dust properties. These elements collectively affect the shape of the continuum spectrum. The introduction of additional parameters to account for systematic biases and correct for dust attenuation further exacerbates this complexity.
@@
By extending the principles behind the absorption line characterisation method, it is conceivable that each pixel in the spectrum can be associated with an “equivalent absorption line”. This approach significantly broadens the applicability of absorption line characterisation, from a mere handful of lines to the entirety of the spectrum. This method integrates the benefits of full-spectrum fitting and absorption line characterisation, not only effectively overcoming the degenerate posed by dust attenuation and flux systematic bias. but also leveraging the rich information embedded in the entire spectrum. Wilkinson et al. (2015, 2017) first developed a code of full spectral fitting (Firefly) based on the premise that dust attenuation and systematic bias primarily influence the large-scale structure of the continuum spectrum, while sparing the small-scale structure of absorption lines. they employed a Fourier-based high-pass filter method (HPF), effectively stripping the observed and model spectra of large-scale features. However, it is crucial to recognize that the HPF approach alone cannot completely eliminate the residual signal from dust attenuation and systematic bias within the small-scale features. To isolate small-scale spectral features from large-scale variations, Li et al. (2020) utilise a moving box average. By analyzing the ratio between these features, they are able to constrain the underlying stellar population. Since dust attenuation impacts both scales uniformly at a given wavelength, the ratio remains unaffected by dust attenuation and systematic bias. However, this method is nonlinear, thereby constraining the parameter space and reducing the operational speed when addressing stellar populations.
@@ 2.1 Invariance of Equivalent Width
The equivalent width ($EW$) is the width of the adjacent continuum that has the same area as is taken up by the absorption or emission line. Mathematically we have:
[…]
where $F^{\prime}$ is the background continuum spectrum. Similarly, we can define the full-spectrum “equivalent width spectrum” as:
@@
The continuum depends on the original spectrum. A well-defined continuum effectively mirrors any global variations in the shape of the original spectrum, encompassing factors like dust attenuation and calibration bias:
$F^{\prime}(\lambda)=F^{\prime}_{init}(\lambda)*P(\lambda).$ (4)
It can be seen that, a well-defined Equivalent Width is invariant in global variations:
$EW(\lambda)=\frac{F(\lambda)}{F^{\prime}(\lambda)}=\frac{F_{init}(\lambda)*P(\lambda)}{F^{\prime}_{init}(\lambda)*P(\lambda)}=\frac{F_{init}(\lambda)}{F^{\prime}_{init}(\lambda)}.$ (5)
@@
$RES(\lambda)=F(\lambda)-F^{\prime}(\lambda)=(F_{init}(\lambda)-F^{\prime}_{init}(\lambda))*P(\lambda).$ (6)
@@ 2.2 Discrete Penalised Least Squares
As mentioned in Section 2.1, a well-defined continuum should reflects any global variations in the shape of the original spectrum. In other words, The initial continuum $F^{\prime}_{init}$ corresponding to the smoothed continuum $F^{\prime}$ derived from Equation 12, adheres to the same Equation 12 when related to the initial continuum $F_{init}$.
@@
$\bm{APF^{\prime}_{init}}=\bm{WF}=\bm{WPF_{init}}$ (13)
@@
$\bm{P^{-1}APF^{\prime}_{init}}=\bm{WF_{init}}$ (14)
@@ 2.4 Global variation and attenuation
Here, $f_{0}$ is a normalization factor to adjust the absolute value of the spectrum, which is unknown. In practice, we relate the global correction function to the V-band ($\lambda_{V}=5500$) to eliminate $f_{0}$:
@@
$-2.5\log\frac{P(\lambda)}{P_{V}}$ $=A(\lambda)-A_{V}-2.5\log\frac{q(\lambda)}{q_{V}}$ (23)
$=-2.5(\log\frac{F(\lambda)}{F_{fit}(\lambda)}-\log\frac{F(\lambda_{V})}{F_{fit}(\lambda_{V})}).$
@@ 2.5 The SEW Extension for pPXF
Figure 1: This figure illustrates the simple mock example spectrum used for our algorithm. The top-right panel shows the initial spectrum, the mock spectrum with attenuation and a signal-to-noise ratio (S/N) of 30, and the smoothed mock spectrum with DPLS method described in Section 2.2 with black, grey and blue line, respectively. The second-right panel shows the “equivalent width spectrum” of initial spectrum, the mock spectrum and the best-fitting spectrum with our method with black, grey and blue dashed lines, respectively. The third-right panel shows the initial spectrum, mock spectrum and best-fit spectrum with black, grey and blue dashed lines, respectively. For comparison, we show the best-fitting results with pPXF+power-law attenuation curve and pPXF+calzetti curve with red and magenta dashed lines. The inset zooms in on a specific wavelength range (3700-4150 Å) to highlight the differences. The bottom-right panel shows different attenuation, with the line types explained in the legend. The four left panels, from top to bottom, show the stellar population weight distributions for the power-law, our algorithm, pPXF+power-law attenuation curve, and pPXF+Calzetti curve, respectively.
@@ 3 Test with mock spectrum
where $E(B-V)$ represents the reddening between the B band and V band, and $\beta$ is the power-law slope.
@@
The top row of Figure 1 displays a spectrum generated through the process described above. In this panel, the left portion illustrates the composition of the stellar population, while the intrinsic spectrum is represented by the black line in the right portion. This mock spectrum is characterised by a luminosity-weighted characteristic logarithmic stellar age $\log\text{age}L$ of $8.5$ and a characteristic metallicity $[Z/Z\odot]$ of $-0.6$, as detailed in the 2nd row of Table 1. To further refine the mock, we apply an attenuation curve with parameters $E(B-V)=0.3$ and $\beta=1.3$. The final mock spectrum, incorporating kinematics, attenuation, and noise with a signal-to-noise ratio (S/N) of $30$, are shown as grey lines.
This mock spectrum is analysed in Section 3.2 using our method to test its performance under controlled conditions. Additionally, to account for potential calibration biases and spectral stitching deviations, we generate two more mock spectra for further testing in Section 3.3. In this paper, we will benchmark the conventional method of pPXF against pPXF-SEW. This is because pPXF is the most accurate full-spectrum fitting method (and the fastest). pPXF is the most accurate full-spectrum fitting method (and the fastest too). In a detailed study, pPXF was found to produce the smallest errors (compared to the known true values) and to be significantly faster than all other full-spectrum fitting codes (Woo et al., 2024). All mocks and corresponding results presented in this section can be reproduced using example files provided with pPXF-SEW.
Table 1: The bestfit $\log age_{L}$ and $[Z/Z_{\odot}]$ of the simple example (Section 3.2), calibration bias example and stitching artefacts example (Section 3.3).
Example Method $\log age_{L}$ $[Z/Z_{\odot}]$
@@
Simple Example SEW 8.51 -0.66
pPXF+mock 8.53 -0.54
pPXF+misfit 8.69 -0.73
Calibration Bias SEW 8.51 -0.66
pPXF+curve 7.98 -0.11
pPXF+poly 8.11 -0.13
stitching artefacts SEW 8.49 -0.64
pPXF+curve 8.78 -1.23
pPXF+poly 8.57 -1.04
3.2 an simple example
@@
The weight fractions are shown in the left portion of the $2^{nd}$ row. The best-fit equivalent width spectrum is plotted in blue in the corresponding right panel. Here, the dashed line indicates the spectrum incorporating emission lines, while the dash-dotted line shows the spectrum without emission lines. A solid grey line marks the intrinsic equivalent width spectrum for comparison. To highlight detailed features, we zoom in on the region from 3700 to 4150 Å (associated with D4000) in the inset box, where abundant emission and absorption lines are visible. The best-fit parameters, $\log age_{L}=8.51$ and $[Z/Z_{\odot}]=-0.66$, are tabulated in Row 3 of Table 1. Although small discrepancies between fitted and mock weight fractions arise from systematic uncertainties, their global distributions show good agreement. Statistical characteristics (including age and metallicity estimates) of these equivalent width spectra match closely with those derived from their intrinsic counterparts. Observed minor deviations likely stem from well-documented age-metallicity degeneracies influencing such measurements.
@@
In the right panel of the third row in Figure 1, we show the full-spectrum reconstruction derived from fitting results using pPXF-SEW. The black line corresponds to the intrinsic mock spectrum, while the blue line shows the SEW method fitted spectrum. For comparison, we additionally performed stellar population synthesis on this mock spectrum using standard pPXF. During this analysis, we applied both our mock attenuation curve (power-law form) and a misfit curve, the Calzetti law curve (Calzetti et al., 2000). The resulting spectra from these two approaches appear as red and magenta dashed lines in the same panel. The derived parameters $\log age_{L}=8.53$ and $[Z/Z_{\odot}]=-0.54$ for the mock attenuation curve compare with $\log age_{L}=8.69$ and $[Z/Z_{\odot}]=-0.73$ for the misfit curve (see Rows 4-5 in Table 1). Corresponding weight fractions appear in left panels 3-4 of Figure 1. An inset magnifies key features between 3700-4150 Å for detailed inspection. The reconstruction of SEW method and pPXF results with mock attenuation curve closely match the intrinsic mock spectrum. Conversely, fits using a misfit curve display marked deviations from ground truth values across multiple spectral features.
@@
Figure 2: Comparison of different spectral processing methods in the presence of flux calibration bias. The top panel shows the full spectrum, including the input spectrum (grey), the mock initial spectrum (black), the EW fit spectrum (bule), and the results from pPXF+mock curve (red) and pPXF+misfit (magenta). The inset zooms in on a specific wavelength range (3700-4150 Å) to highlight the differences. The bottom panel displays the attenuation curve, showing the mock attenuation (dashed line), the calibration bias(dotted line) , the mock attenuation with bias (solid line), the EW fit results (blue dots and line), and the results from pPXF + mock curve (red) and pPXF+misfit (magenta).
@@
3.3 mock examples with flux bias
@@
For flux calibration bias, we introduce periodic perturbations through sinusoidal modulation (1000 Å wavelength periodicity, 0.05 mag amplitude; see black dashed curve in Figure 2, lower panel). Spectral stitching artefacts are modelled via a localized 0.1 mag flux depression –applied exclusively to the blue segment ($\lambda<5100$) at the merging boundary (black dashed line in Figure 3, lower panel).
Figures 2 and 3 present our recovery results using standardized panel layouts: upper panels compare reconstructed unattenuated spectra (colour-coded by solution type), while lower panels display corresponding recovered attenuation curves. Quantitative stellar population parameters ($\log age_{L}$ and $[Z/Z_{\odot}]$) are systematically tabulated in Table 1.
@@
Conventional SPS methods lacking explicit bias correction mechanisms (pPXF+curve case) exhibit significant deviations from true mock values, as demonstrated by discrepancies in recovered $\log age_{L}$ and $[Z/Z_{\odot}]$ parameters, along with mismatches in reconstructed intrinsic spectra and evolutionary curves (red lines in Figures 2 and 3) when compared with SEW results and ground truth simulations.
In contrast, parametric bias calibration approaches (pPXF+poly case) fail to achieve satisfactory bias correction even when employing tenth-order polynomials ($\log age_{L}$ and $[Z/Z_{\odot}]$ parameters of pPXF+poly in Table 1 and magenta lines in Figures 2 and 3). Furthermore, this nonlinear optimization process substantially increases computational costs during SPS fitting - in our test case, the pPXF+poly implementation exceeded the runtime of the SEW method.

### v2

@@ 1 Introduction
In Galaxy spectra analysis, two primary stellar population synthesis (SPS) methods are commonly employed: the full-spectrum fitting and the absorption line characterisation method.
In full-spectrum fitting, observed spectra are modelled through linear combinations of SSPs, each characterised by distinct age and metallicity parameters. These SSP templates incorporate integration over an initial mass function (IMF) and stellar evolutionary tracks. The fitting process adjusts the weights of SSPs, along with other parameters, to minimise residuals between the model and data:
$F(\lambda)$ $=F_{\text{intr}}(\lambda)*P(\lambda),$ (1a)
$F_{\text{intr}}(\lambda)$ $=\sum_{j}x_{j}f_{\text{temp},j}(\lambda),$ (1b)
@@
Here, $F_{\text{intr}}(\lambda)$ represents the intrinsic, unattenuated galactic spectrum comprising SSP and emission-line templates $f_{\text{temp}}^{(j)}(\lambda)$ with corresponding weights $x_{j}$. The observed spectrum $F(\lambda)$ emerges through modification by the global variations $P(\lambda)$, which accounts for both intrinsic (self-) and foreground Galactic extinction $A(\lambda)$, combined with instrumental calibration deviations $q(\lambda)$. This inverse modelling approach enables the recovery of intrinsic galactic properties from observational data through spectral decomposition. Several established spectral synthesis codes implement this methodology, including pPXF (Cappellari & Emsellem, 2004; Cappellari, 2017), STARLIGHT (Cid Fernandes et al., 2005), STECKMAP (Ocvirk et al., 2006a, b), and VESPA (Tojeiro et al., 2007). However, the accuracy of the fitting results is significantly influenced by several interrelated factors, including the choice of the theoretical dust attenuation curve, the precision of spectral calibrations, and the inherent degeneracy among age, metallicity, and dust properties. These elements collectively affect the shape of the continuum spectrum. The introduction of additional parameters to account for systematic biases and correct for dust attenuation further exacerbates this complexity.
@@
By extending the principles behind the absorption line characterisation method, it is conceivable that each pixel (whether or not absorption line features exhibit) in the spectrum can be associated with an “equivalent absorption line”. This approach significantly broadens the applicability of absorption line characterisation, from a mere handful of lines to the entirety of the spectrum. This method integrates the benefits of full-spectrum fitting and absorption line characterisation, not only effectively overcoming the degenerate posed by dust attenuation and flux systematic bias. but also leveraging the rich information embedded in the entire spectrum. Wilkinson et al. (2015, 2017) first developed a code of full spectral fitting (Firefly) based on the premise that dust attenuation and systematic bias primarily influence the large-scale structure of the continuum spectrum, while sparing the small-scale structure of absorption lines. they employed a Fourier-based high-pass filter method (HPF), effectively stripping the observed and model spectra of large-scale features. However, it is crucial to recognize that the HPF approach alone cannot completely eliminate the residual signal from dust attenuation and systematic bias within the small-scale features. To isolate small-scale spectral features from large-scale variations, Li et al. (2020) utilise a moving box average. By analyzing the ratio between these features, they are able to constrain the underlying stellar population. Since dust attenuation impacts both scales uniformly at a given wavelength, the ratio remains unaffected by dust attenuation and systematic bias. However, this method is nonlinear, thereby constraining the parameter space and reducing the operational speed when addressing stellar populations.
@@ 2.1 Invariance of Equivalent Width
The equivalent width ($EW$) is the width of the adjacent continuum $F^{\prime}(\lambda)$ that has the same area as is taken up by the absorption or emission line in original spectrum $F(\lambda)$. Mathematically we have:
$EW=\int\frac{F(\lambda)-F^{\prime}(\lambda)}{F^{\prime}(\lambda)}d\lambda,$ (2)
where $F^{\prime}$ is the background smoothed continuum spectrum [^2] The superscript ’ in the following text indicates the smoothing continuum of the corresponding spectrum.. Similarly, we can define the full-spectrum “equivalent width spectrum” as:
@@
The continuum $F^{\prime}(\lambda)$ depends on the original spectrum $F(\lambda)$. A well-defined continuum effectively mirrors any global variations $P(\lambda)$ in the shape of the original spectrum, encompassing factors like dust attenuation and calibration bias:
$F^{\prime}(\lambda)=F^{\prime}_{intr}(\lambda)*P(\lambda),$ (4)
where the $F^{\prime}_{intr}(\lambda)$ is the continuum of intrinsic original spectrum $F_{intr}(\lambda)$ that this equation corresponds to Equation 1a.
It can be seen that, a well-defined Equivalent Width is invariant in global variations $P(\lambda)$:
$EW(\lambda)=\frac{F(\lambda)}{F^{\prime}(\lambda)}=\frac{F_{intr}(\lambda)*P(\lambda)}{F^{\prime}_{intr}(\lambda)*P(\lambda)}=\frac{F_{intr}(\lambda)}{F^{\prime}_{intr}(\lambda)}.$ (5)
@@
$RES(\lambda)=F(\lambda)-F^{\prime}(\lambda)=(F_{intr}(\lambda)-F^{\prime}_{intr}(\lambda))*P(\lambda).$ (6)
@@ 2.2 Discrete Penalised Least Squares
As mentioned in Section 2.1, a well-defined continuum should reflects any global variations in the shape of the original spectrum. In other words, The intrinsic continuum $F^{\prime}_{intr}$ corresponding to the smoothed continuum $F^{\prime}$ derived from Equation 12, adheres to the same Equation 12 when related to the intrinsic continuum $F_{intr}$.
@@
$\bm{APF^{\prime}_{intr}}=\bm{WF}=\bm{WPF_{intr}}$ (13)
@@
$\bm{P^{-1}APF^{\prime}_{intr}}=\bm{WF_{intr}}$ (14)
@@ 2.4 Global variation and attenuation
Here, $f_{0}$ is a normalization factor to adjust the absolute value of the spectrum, which is unknown. In practice, we relate the global correction function to the V-band ($\lambda_{V}=5500$ Å) to eliminate $f_{0}$:
@@
[…]
@@ 2.5 The SEW Extension for pPXF
Figure 1: This figure illustrates the simple mock example spectrum used for our algorithm. Top-right: Spectral components showing the intrinsic spectrum (black) generated with the top-left SSPs distribution, attenuated mock spectrum with S/N=30 (grey, with the attenuation curve shown as black dashed lines in the bottom-right panel), and DPLS-smoothed spectrum (blue; Section 2.2). Second-right: Equivalent width spectrum comparison between intrinsic (black), attenuated mock (grey), and SEW best-fit (blue dashed) with corresponding SSP weights in second-left panels. Third-right: Spectral reconstructions (combining Equation 1b and the corresponding SSPs weights of each method) comparing intrinsic (black), attenuated mock (grey), SEW best-fit (blue dashed; second-left weights), pPXF+power-law (red dashed; third-left weights), and pPXF+misfit (magenta dashed; fourth-left weights). The inset highlights 3700-4150 Å features. Bottom-right: Comparison of attenuation curves derived from different methods. The attenuation curve of SEW is derived by applying Equation 23 to the best-fit spectrum of SEW and mock spectrum (blue and grey lines in third-right panel respectively), while the pPXF attenuation curve directly represents the parametric form derived from its fitting results. Line styles correspond to methodologies as detailed in the legend.
@@ 3.1 Generating mock spectra
where E(BV)E(B-V) represents the reddening between the B band and V band, and $\beta$ is the power-law slope.
@@
The top row of Figure 1 displays a spectrum generated through the process described above. In this panel, the left portion illustrates the composition of the stellar population, while the intrinsic spectrum is represented by the black line in the right portion. This mock spectrum is characterised by a luminosity-weighted characteristic logarithmic stellar age $\log\text{age}_{L}/yr$ of $8.5$ and a characteristic metallicity $Z/Z\odot]$ of 0.6-0.6, as detailed in the 2nd row of Table 1. To further refine the mock, we apply an attenuation curve with parameters E(BV)=0.3E(B-V)=0.3 and $\beta=1.3$. The final mock spectrum, incorporating kinematics, attenuation, and noise with a signal-to-noise ratio (S/N) of $30$, are shown as grey lines.
This mock spectrum is analysed in Section 3.2 using our method to test its performance under controlled conditions. Additionally, to account for potential calibration biases and spectral stitching deviations, we generate two more mock spectra for further testing in Section 3.3. In this paper, we focus our comparative analysis on benchmarking pPXF-SEW against the conventional pPXF method to evaluate the robustness against extinction and calibration biases of SEW method. Readers interested in broader comparisons with other stellar population synthesis algorithms may consult Woo et al. (2024) for comprehensive testing results. All mocks and corresponding results presented in this section can be reproduced using example files provided with pPXF-SEW.
Table 1: The bestfit $\log\text{age}_{L}/yr$ and $Z/Z_{\odot}]$ of the simple example (Section 3.2), calibration bias example and stitching artefacts example (Section 3.3). Here, the errors are estimated by non-negative least squares of the covariance matrix.
Example Method $\log\text{age}_{L}/yr$ $Z/Z_{\odot}]$
@@
Simple Example SEW $8.5\pm 0.18$ $0.65\pm 0.23$
pPXF+mock $8.53\pm 0.18$ $0.54\pm 0.57$
pPXF+misfit $8.69\pm 0.10$ $0.73\pm 0.15$
Calibration Bias SEW $8.51\pm 0.17$ $0.66\pm 0.18$
pPXF+curve $7.98\pm 0.04$ $0.11\pm 0.05$
pPXF+poly $8.11\pm 0.02$ $0.13\pm 0.03$
stitching artefacts SEW $8.49\pm 0.14$ $0.64\pm 0.18$
pPXF+curve $8.78\pm 0.16$ $1.23\pm 0.16$
pPXF+poly $8.57\pm 0.04$ $1.04\pm 0.07$
3.2 Bias-free attenuation with SEW
@@
The weight fractions are shown in the left portion of the $2^{nd}$ row. The best-fit equivalent width spectrum corresponding to the weight is plotted in blue in the corresponding right panel. Here, the dashed line indicates the spectrum incorporating emission lines, while the dash-dotted line shows the spectrum without emission lines. A solid grey line marks the intrinsic equivalent width spectrum for comparison. To highlight detailed features, we zoom in on the region from 3700 to 4150 Å (associated with D4000) in the inset box, where abundant emission and absorption lines are visible. The best-fit parameters, $\log\text{age}_{L}/yr=8.51$ and $Z/Z_{\odot}]=-0.66$, are tabulated in Row 3 of Table 1. The linearisation is what allows us to obtain the corresponding weights for each SSP. Although small discrepancies between fitted and mock weight fractions arise from numerical effects and systematic uncertainties, their global distributions and spectral profiles show good agreement. Statistical characteristics (including age and metallicity estimates) of these equivalent width spectra match closely with those derived from their intrinsic counterparts. Observed minor deviations likely stem from well-documented age-metallicity degeneracies influencing such measurements.
@@
In the right panel of the third row in Figure 1, we show the full-spectrum reconstruction derived from fitting results using pPXF-SEW with Equation 1b. The black line corresponds to the intrinsic mock spectrum, while the blue line shows the SEW method fitted spectrum. For comparison, we additionally performed stellar population synthesis on this mock spectrum using standard pPXF. During this analysis, we applied both our mock attenuation curve (power-law form) and a misfit curve, the Calzetti law curve (Calzetti et al., 2000). The resulting spectra from these two approaches appear as red and magenta dashed lines in the same panel. The derived parameters $\log\text{age}_{L}/yr=8.53$ and $Z/Z_{\odot}]=-0.54$ for the mock attenuation curve compare with $\log\text{age}_{L}/yr=8.69$ and $Z/Z_{\odot}]=-0.73$ for the misfit curve (see Rows 4-5 in Table 1). Corresponding weight fractions appear in left panels 3-4 of Figure 1. An inset magnifies key features between 3700-4150 Å for detailed inspection. The reconstruction of SEW method and pPXF results with mock attenuation curve closely match the intrinsic mock spectrum. Conversely, fits using a misfit curve display marked deviations from ground truth values across multiple spectral features.
@@
Figure 2: Comparison of different spectral processing methods in the presence of flux calibration bias. The top panel shows the full spectrum, including the input spectrum (grey), the mock intrinsic spectrum (black), the EW fit spectrum (bule), and the results from pPXF+mock curve (red) and pPXF+misfit (magenta). The inset zooms in on a specific wavelength range (3700-4150 Å) to highlight the differences. The bottom panel displays the attenuation curve, showing the mock attenuation (dashed line), the calibration bias(dotted line) , the mock attenuation with bias (solid line), the EW fit results (blue dots and line), and the results from pPXF + mock curve (red) and pPXF+misfit (magenta).
@@
3.3 Flux bias robustness of SEW
@@
For flux calibration bias, we introduce periodic perturbations through sinusoidal modulation (1000 Å wavelength periodicity, 0.05 mag amplitude; see black dashed curve in Figure 2, lower panel). Spectral stitching artefacts are modelled via a localized 0.1 mag flux depression –applied exclusively to the blue segment ($\lambda<5100$ Å) at the merging boundary (black dashed line in Figure 3, lower panel). The introduced bias settings are informed by potential artifacts encountered in real observational scenarios: Periodic gain variations caused by circuit design limitations may produce the flux calibration bias described in our setup, while calibration discrepancies between red and blue CCD detectors could generate spectral stitching artifacts. To accentuate the systematic effects of calibration biases, we deliberately configured these artificial perturbations with heightened amplitude and non-smooth variations.
Figures 2 and 3 present our recovery results using standardized panel layouts: upper panels compare reconstructed unattenuated spectra (colour-coded by solution type), while lower panels display corresponding recovered attenuation curves. Quantitative stellar population parameters ($\log\text{age}_{L}/yr$ and $Z/Z_{\odot}]$) are systematically tabulated in Table 1.
@@
Conventional SPS methods lacking explicit bias correction mechanisms (pPXF+curve case) exhibit significant deviations from true mock values, as demonstrated by discrepancies in recovered $\log\text{age}_{L}/yr$ and $Z/Z_{\odot}]$ parameters, along with mismatches in reconstructed intrinsic spectra and evolutionary curves (red lines in Figures 2 and 3) when compared with SEW results and ground truth simulations.
In contrast, parametric bias calibration approaches (pPXF+poly case) fail to achieve satisfactory bias correction even when employing tenth-order polynomials ($\log\text{age}_{L}/yr$ and $Z/Z_{\odot}]$ parameters of pPXF+poly in Table 1 and magenta lines in Figures 2 and 3). Furthermore, this nonlinear optimization process substantially increases computational costs during SPS fitting - in our test case, the pPXF+poly implementation exceeded the runtime of the SEW method.